# Judging the Judges: Can Large Vision-Language Models Fairly Evaluate Chart Comprehension and Reasoning?

### v1

@@ Abstract
Md Tahmid Rahman Laskar‡,thanks:   Contact Emails: {tahmedge,enamulh,jhuang}@yorku.ca, Mohammed Saidul Islam‡,thanks:   Equal Contributions., Ridwan Mahbub‡,22footnotemark: 2, Ahmed Masry, Mizanur Rahman, Md Amran Hossen Bhuiyan, Mir Tafseer Nayeem§, Shafiq Joty, Enamul Hoque‡,11footnotemark: 1, Jimmy Xiangji Huang‡,11footnotemark: 1 York University, §University of Alberta, Salesforce AI Research
@@ 1 Introduction
To this end, this paper aims to investigate whether open-source smaller LVLMs (e.g., less than 10B parameters) can effectively evaluate answers about charts – assessing correctness, relevance, and other qualities – similarly to a human or a powerful LLM like GPT-4 OpenAI et al. (2023)? For this purpose, we conduct one of the first comprehensive evaluations of open-source LVLMs as evaluators on various chart benchmarks, consisting of diverse tasks like chart captioning and question answering. We focus on open-source, smaller VLMs (up to 10B parameters) to simulate realistic deployment scenarios where cost-effective or private models are preferred over large closed models. By benchmarking these models against high-quality reference judgments generated by closed-source LLMs like GPT-4 or 70B open-source LLM-Judge like LLaVA-Critic Xiong et al. (2024), we aim to uncover to what extent current open models can serve as reliable judges, and when they fail.
@@
(1) We establish an evaluation framework for chart comprehension using “LVLM-as-a-Judge”, with clear rubrics for pairwise and pointwise assessments over 100K judgments generated by GPT-4o and LLaVA-Critic-70B. Additionally, we introduce a new benchmark to assess the instruction-following abilities of LVLMs in Chart related tasks.
(2) We evaluate a wide range of open-source multimodal LLMs as judges – 13 models ranging from 2B to 9B parameters – and analyze their performance against LLM-annotated (GPT-4 and LLaVA-Critic) and human-annotated reference judgments, across diverse chart benchmarks (OpenCQA and VisText) on answers generated by different LLMs to create challenging evaluation scenarios.
(3) We provide an in-depth analysis of the judges’ strengths and weaknesses, revealing issues like position bias and length bias, and discuss which models achieve substantially higher agreement with reference judgments, and which ones fail.
In addition, our code, judgment data, and our proposed instruction-following evaluation benchmark is released here: https://github.com/tahmedge/chart_lvlm_judge.
@@ 4 Experiments
In this section, we present the experimental results based on evaluating 13 LVLMs as judges across OpenCQA, VisText, and our proposed Chart-Instruct-Eval. The evaluation considers both pairwise and pointwise assessments, focusing on factual correctness, informativeness, relevance, positional bias, length bias, and instruction-following accuracy. We parse the LVLM-judge predicted judgments from their corresponding responses using a parsing script Laskar et al. (2023, 2024a, 2024b). Note that we ran all our experiments using 1 A100 GPU with all the decoding parameters being set to the default values in HuggingFace Wolf et al. (2020). Below, we demonstrate our findings:
@@ 4.2 Pointwise Evaluation Results
ii. Lower-performing models: Similar to the pairwise scenario, PaliGemma-3B and ChartGemma-3B again produced irrelevant outputs resulting in the highest error distances (5.0). Moreover, despite being the largest model in our evaluation, the Idefics-9B-Instruct model performs quite poorly with very high error distance (on average, above 3).
@@ 4.3 Instruction and Format Adherence
We also assess the LVLM judges on their ability to maintain a standardized response format and whether they can evaluate the instruction following capabilities of other models. Based on the results in Table 2, we find that all 7B models achieve more than 90% format following capability. Smaller LVLMs like Qwen-2B and Phi-3.8B also achieve around 80% format adherence. In terms of instruction following capability evaluation, mPLUG-Owl3-7B (93.5%) and Qwen2-VL-7B-Instruct (87.0%) achieve the top two results. Surprisingly, the LLaVA-Critic-7B model achieves only 45.5% accuracy in this task. This may indicate that the training data of the LLaVA-Critic-7B model may not contain such data, leading to a quite poor generalization in this dataset. Moreover, PaliGemma-3B and ChartGemma-3B failed to follow the format requirements at all (also unable to evaluate instruction following capability). Finally, the Idefics-9B-Instruct model, even with 9B parameters achieves poor instruction and format following accuracy.
@@ 5 Conclusion and Future Work
In this paper, we conducted a comprehensive evaluation of open-source LVLMs as automatic judges for chart comprehension and reasoning tasks. Our analyses revealed that while some open-source LVLMs (e.g., 7B models like LLaVA-Critic, Qwen2-VL, InternLM, and LLaVA-Next) can achieve judgment accuracy (with lower error rates) which is comparable to state-of-the-art closed-source models like GPT-4 or larger open-source models like LLaVA-Critic-70B; other models, such as ChartGemma and PaliGemma, struggle significantly, highlighting variability in their reliability. Despite the promising results of various models, issues like bias and lack of instruction following capability still persist. Therefore, future work should focus on mitigating biases, improving instruction following evaluation capability, alongside ensuring consistency across diverse evaluation criteria by developing a multimodal LLM judge tailored for chart model evaluation.
@@ Acknowledgments
This research was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada, the York Research Chairs (YRC) program, and Compute Canada.
@@ A.1 Regarding Model and Dataset Selection
With the rapid development of numerous LVLMs recently, our evaluation may not cover all LVLMs Grattafiori et al. (2024). Nonetheless, we tried to use as many state-of-the-art LVLMs as possible (we mainly focused on popular LVLMs released by the beginning of 2025, with sizes less than 10B parameters). While some LVLMs that are proposed at the time of writing this paper and not used in this study may yield different results Bai et al. (2025), the findings from our study will still offer substantial insights into building better-automated evaluators and chart-specific LLM judges in the future. Although there are many other chart benchmarks currently available, not all benchmarks (e.g., ChartQA Masry et al. (2022), fact-checking over chart images Akhtar et al. (2023, 2024)) require qualitative evaluation. Therefore we selected OpenCQA and VisText since qualitative evaluation is often required in these datasets Islam et al. (2024). Moreover, we prioritize VisText over other chart captioning datasets since it contains both L1 and L2/L3 type captions.

### v2

@@ Abstract
Md Tahmid Rahman Laskar‡,thanks:   Contact Emails: {tahmedge,enamulh,jhuang}@yorku.ca, Mohammed Saidul Islam‡,thanks:   Equal Contributions., Ridwan Mahbub‡,22footnotemark: 2, Ahmed Masry, Mizanur Rahman, Md Amran Hossen Bhuiyan, Mir Tafseer Nayeem§, Shafiq Joty¶,\$, Enamul Hoque‡,11footnotemark: 1, Jimmy Xiangji Huang‡,11footnotemark: 1 York University, §University of Alberta, Nanyang Technological University, \$Salesforce AI Research
@@ 1 Introduction
To this end, this paper aims to investigate whether open-source smaller LVLMs (e.g., less than 10B parameters) can effectively evaluate answers about charts—assessing correctness, relevance, and other qualities—similarly to a human or a powerful LLM like GPT-4 OpenAI et al. (2023). For this purpose, we conduct one of the first comprehensive evaluations of open-source LVLMs as evaluators on various chart benchmarks, consisting of diverse tasks like chart captioning and question answering. We focus on open-source, smaller VLMs (up to 10B parameters) to simulate realistic deployment scenarios where cost-effective or private models are preferred over large closed models. By benchmarking these models against high-quality reference judgments generated by closed-source LLMs like GPT-4 or 70B open-source LLM-Judge like LLaVA-Critic Xiong et al. (2024), we aim to uncover to what extent current open models can serve as reliable judges, and when they fail.
@@
  1. 1.
    We establish an evaluation framework for chart comprehension using “LVLM-as-a-Judge”, with clear rubrics for pairwise and pointwise assessments over 100K judgments generated by GPT-4o and LLaVA-Critic-70B. Additionally, we introduce a new benchmark to assess the instruction-following abilities of LVLMs in chart-related tasks.
  2. 2.
    We evaluate a wide range of open-source multimodal LLMs as judges – 13 models ranging from 2B to 9B parameters – and analyze their performance against LLM-annotated (GPT-4 and LLaVA-Critic) and human-annotated reference judgments, across diverse chart benchmarks (OpenCQA and VisText) on answers generated by different LLMs to create challenging evaluation scenarios.
  3. 3.
    We provide an in-depth analysis of the judges’ strengths and weaknesses, revealing issues like position bias and length bias, and discuss which models achieve substantially higher agreement with reference judgments, and which ones fail.
In addition, our code, judgment data, and our proposed instruction-following evaluation benchmark is released here: https://github.com/tahmedge/chart_lvlm_judge
@@ 4 Experiments
In this section, we present the experimental results based on evaluating 13 LVLMs as judges across OpenCQA, VisText, and our proposed Chart-Instruct-Eval. The evaluation considers both pairwise and pointwise assessments, focusing on factual correctness, informativeness, relevance, positional bias, length bias, and instruction-following accuracy. We parse the LVLM-judge predicted judgments from their corresponding JSON-formatted responses using a parsing script Laskar et al. (2023, 2024a, 2024b). If the parsing script cannot properly parse the judgment from the response, we consider the LLM-generated answer as wrong for the pairwise case and error distance of 5 for the pointwise case. Note that we ran all our experiments using 1 A100 GPU with all the decoding parameters being set to the default values in HuggingFace Wolf et al. (2020). Below, we demonstrate our findings:
@@ 4.2 Pointwise Evaluation Results
ii. Lower-performing models: Similar to the pairwise scenario, PaliGemma-3B and ChartGemma-3B again produced irrelevant outputs resulting in the highest error distances (5.0). Moreover, despite being the largest model in our evaluation, the Idefics-9B-Instruct model performs quite poorly with a high error (on average, above 3).
@@ 4.3 Instruction and Format Adherence
We also assess the LVLM judges on their ability to maintain a standardized response format and whether they can evaluate the instruction following capabilities of other models. Based on the results in Table 2, we find that all 7B models achieve more than 90% format following capability. Smaller LVLMs like Qwen-2B and Phi-3.8B also achieve around 80% format adherence.
In terms of instruction following capability evaluation, we find that many LVLMs that could properly follow the format following requirement in their generated judgments for pairwise (§4.1) and pointwise (§4.2) evaluations, surprisingly generate the response in the wrong format in this evaluation. This makes our original parsing script penalize most of the LVLM-generated judgments as wrong. Therefore, we rewrite the parsing script to make it more flexible in terms of format adherence of the LVLM judge, since for this evaluation, our focus was to evaluate whether LVLM judges can properly assess instruction-following capabilities of different models in downstream chart-related tasks. Therefore, format adherence and other capabilities of the LVLM judges were not the focus of this evaluation. Our experiments reveal that mPLUG-Owl3-7B (93.5%) and Qwen2-VL-7B-Instruct (87.0%) achieve the top two results in terms of evaluating the instruction-following capability of different LVLM generated responses. Surprisingly, the LLaVA-Critic-7B model achieves only 45.5% accuracy in this task. This may indicate that the training data of the LLaVA-Critic-7B model may not contain such data, leading to a quite poor generalization in this dataset.
Moreover, PaliGemma-3B and ChartGemma-3B fail to follow the format requirements at all, and also unable to evaluate instruction following capability. Finally, the Idefics-9B-Instruct model, even with 9B parameters, achieves poor instruction and format following accuracy.
@@ 5 Conclusion and Future Work
In this paper, we conducted a comprehensive evaluation of open-source LVLMs as automatic judges for chart comprehension and reasoning tasks. Our analyses revealed that while some open-source LVLMs (e.g., 7B models like LLaVA-Critic, Qwen2-VL, InternLM, and LLaVA-Next) can achieve judgment accuracy (with lower error rates) that is comparable to state-of-the-art closed-source models like GPT-4 or larger open-source models like LLaVA-Critic-70B; other models, such as ChartGemma and PaliGemma, struggle significantly, highlighting variability in their reliability. Despite the promising results of various models, issues like bias and lack of instruction following capability still persist. Therefore, future work should focus on mitigating biases, improving instruction following evaluation capability, alongside ensuring consistency across diverse evaluation criteria by developing a multimodal LLM judge using more recent models Bai et al. (2025).
@@ Acknowledgments
This research is supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada, the York Research Chairs (YRC) program, and Compute Canada. This paper is accepted at the ACL 2025 Industry Track and we thank all the anonymous reviewers for their excellent review comments. We would also like to thank Mehrad Shahmohammadi for his help in this research.
@@ A.1 Regarding Model and Dataset Selection
We selected popular LVLMs that were released by early 2025, with sizes less than 10B parameters. Although there are many other chart benchmarks currently available Huang et al. (2024), we selected OpenCQA and VisText since qualitative evaluation is often required in these datasets Islam et al. (2024).